OpenStudy (anonymous):

Evaluate the function for x = 8. f(x)=x^2/16-2x A. –15 B. 20 C. –12 D. –24

OpenStudy (anonymous):

plug in 8 for x and solve

OpenStudy (anonymous):

\[f(x)= \frac{ x^2 }{ 16-2x } \] \[f(x)= \frac{ (8)^2 }{ 16-2(8) } \] \[f(x)= \frac{ 64 }{ 16-16 } \] NO SOLUTION

OpenStudy (anonymous):

Did you write it correct?

OpenStudy (anonymous):